# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 6.2. Structure constants under flops
As we have $\overline{\mathcal{X}}=\overline{\mathcal{X}}^{\prime}$, it follows that $\mathcal{X}$ and $\mathcal{X}^{\prime}$ are related by a small modification over $\overline{\mathcal{X}}$, that is, by a composition of elementary modifications, which, by Lemma 6.1, are either M1 or M2 flops. Given three quasi-projective open Kulikov degenerations $\mathcal{X}$, $\mathcal{X}^{\prime}$, and $\mathcal{X}^{\prime\prime}$ with $\overline{\mathcal{X}}=\overline{\mathcal{X}}^{\prime}=\overline{\mathcal{X}}^{\prime\prime}$, we have $\varphi^{\prime\prime}-\varphi=(\varphi^{\prime\prime}-\varphi^{\prime})+(\varphi^{\prime}-\varphi)$, and so it is enough to prove Theorem 6.6 for $\mathcal{X}$ and $\mathcal{X}^{\prime}$ related either by an M1 or an M2 flop.
If $\mathcal{X}^{\prime}$ and $\mathcal{X}$ are related by an M1 flop, then $\mathcal{X}_{0}$ and $\mathcal{X}_{0}^{\prime}$ have the same double curves. In particular, we have $\varphi=\varphi^{\prime}$, and in this case Theorem 6.6 follows from Theorem 6.2. If $\mathcal{X}^{\prime}$ and $\mathcal{X}$ are related by an M2 flop, then we have $\varphi^{\prime}-\varphi=\varphi_{E}[C]$ by Lemma 6.5, and so in this case Theorem 6.6 follows from Theorem 6.4. ∎
@@ 6.3. The finiteness property
In this section, we prove a finiteness result for the punctured Gromov–Witten invariants $N_{pqr}^{\beta}(\mathcal{X})$, which implies that the mirror algebra of a quasi-projective open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$ can be defined over $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$, and not just over the Artinian rings $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]/I$. We first introduce the notion of a good divisor in Definition 6.8 and prove the finiteness result under the assumption of existence of a good divisor in Theorem 6.11. Then, we show in Lemmas 6.12-6.13 that a good divisor always exist after flops, and we deduce from this the general finiteness result in Theorem 6.16.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. For every divisor $F$ on $\mathcal{X}$ supported on $\mathcal{X}_{0}$, and for every $p,q,r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$, we have
@@
    The divisor $-F$ is nef, and for any compact irreducible curve $C\subset\mathcal{X}_{0}$ we have $-F\cdot C\geq 1$, except for finitely many disjoint curves $C$ which are either internal exceptional curves in an irreducible component of $\mathcal{X}_{0}$, or irreducible components $D_{ij}=X^{i}\cap X^{j}$ of the double locus of $\mathcal{X}_{0}$, satisfying $D_{ij}|_{X^{i}}^{2}=D_{ij}|_{X^{j}}^{2}=-1$.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and $F$ a good effective divisor on $\mathcal{X}$. By Definition 6.8 ii) of a good divisor, the curves $C$ with $F\cdot C=0$ are exactly the curves contracted by a birational morphism, and so they span a face $G$ of the cone $NE(\mathcal{X}/\overline{\mathcal{X}})\otimes\mathbb{R}$. We denote by $J\subset NE(\mathcal{X}/\overline{\mathcal{X}})$ the corresponding monoid ideal consisting of curve classes $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $\beta\notin G$. By Definition 6.8 i) of a good divisor, the curve classes $\beta\in J$ are exactly the curve classes $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $-F\cdot\beta\geq 1$.
@@
where $v$ and $v^{\prime}$ are the vertices adjacent to $e$, $(E_{i})_{i\in I_{v,e}}$ (resp. $(E_{j})_{j\in I_{v^{\prime},e}}$) is the set of internal exceptional curves $E_{i}$ in $X^{v}$ (resp . $X^{v^{\prime}}$) intersecting $X^{e}$, and where $m_{v,e}$ (resp.) is the integral primitive direction of $e$ pointing towards $v$ (resp . $v^{\prime}$).
@@
We have to consider $\mathbb{A}^{1}$-curves $C$ of class $[C]\notin J$, that is, $[C]\in G$. As $C$ is connected, and the curves in Definition 6.8 i) spanning $G$ are disjoint and rigid, it follows that either $C$ is a multiple of an internal exceptional curve or a multiple of a double curve. The formal neighborhood of a double curve is toric, whereas the tropicalization of an $\mathbb{A}^{1}$-curve has a unique leg. Hence, it follows by the toric balancing condition for punctured curves that $C$ cannot be a multiple of a double curve. Thus, $C$ is a multiple of an internal exceptional curve $E_{i}$ of an irreducible component $X^{v}$ of $\mathcal{X}_{0}$, and intersecting a component $X^{e}\subset X^{v}$ of the double locus of $\mathcal{X}_{0}$. Denote by $f_{W,E_{i}}$ the contribution to $f_{W}$ of the multiple covers of $E_{i}$. The formal neighborhood of $E_{i}$ in the log scheme $\mathcal{X}_{0}^{\dagger}$ is isomorphic to the formal neighborhood of the exceptional curves appearing in the proof of [18, Lemma 4.21], and so it follows from [18, Lemma 4.21] that we have
@@
We prove the result by induction on $k$. For $k=0$, the result follows from Lemma 6.9. For the induction step, let us assume that the result holds for $k-1$, and show that it holds for $k$. For this, first note that every wall $(W,f_{W})$ of $\mathfrak{D}_{\mathcal{X}}$ with $f_{W}\neq 1\mod J^{k}$ and $f_{W}=1\mod J^{k-1}$ has an initial point $x\in P$. By consistency of $\mathfrak{D}_{\mathcal{X}}$, either $x\in P_{\mathbb{Z}}$, or $x$ is an intersection point of walls which are non-trivial modulo $J^{k-1}$. By the induction hypothesis, there are finitely many walls non-trivial modulo $J^{k-1}$, and so finitely many possible points $x$ which are initial points for a wall non-trivial modulo $J^{k}$. Hence, it is enough to prove that for such a point $x$, there are finitely many walls $(W,f_{W})$ with initial point $x$ such that $f_{W}\neq 1\mod J^{k}$, and that for every such a wall, $f_{W}\mod J^{k}$ is a polynomial.
@@
If $x$ is contained in the interior of a 2-dimensional face of $\mathscr{P}$, then $\mathfrak{D}_{\mathcal{X}}$ is given locally around $x$ by a consistent scattering diagram in $\mathbb{R}^{2}$. Moreover, all the walls passing through $x$ are trivial modulo $J$ by Lemma 6.9, and so the walls non-trivial modulo $J^{k}$ are uniquely determined by consistency from the walls non-trivial modulo $J^{k-1}$, and so the finiteness and polynomiality of the walls non-trivial modulo $J^{k}$ follows from the induction hypothesis and [39, Theorem 6.38].
@@
If $x$ is contained in the interior of an edge of $\mathscr{P}$, then $\mathfrak{D}_{\mathcal{X}}$ is still given locally around $x$ by a consistent scattering diagram in $\mathbb{R}^{2}$. The only difference with the previous case is that some walls passing through $x$ might be trivial modulo $J$. However, these walls are of the explicit form given by Lemma 6.9, and so the result still holds by [39, Proposition 6.47].
@@
It remains to consider the case where $x=v$ is a vertex of $\mathscr{P}$, corresponding to an irreducible component $X^{v}$ of $\mathcal{X}_{0}$. It follows from [40, Theorem 8.15] that the punctured Gromov–Witten invariants contributing to the walls $(W,f_{W})$ with initial point $x=v$ and non-trivial modulo $J^{k}$ can be expressed explicitly in terms of punctured Gromov–Witten invariants $N_{\tau}^{\beta}$ of $(X^{v},\partial X_{v})$ and of the walls non-trivial modulo $J^{k-1}$ passing through $v$. More precisely, it follows from [40, Lemma 8.9] that, up to applying corner blow-ups to $(X^{v},\partial X^{v})$, the invariants $N_{\tau}^{\beta}$ are counts of curves $C$ in $X^{v}$ intersecting $\partial X^{v}$ at fixed smooth points $x_{1},\dots,x_{q}$, with multiplicities $m_{1},\dots,m_{q}$, and at a smooth non-fixed point $x_{\mathrm{out}}$ with multiplicity $m_{\mathrm{out}}$. If $q\neq 1$, then, as in the proof of [46, Lemma 4.1], given the points $x_{1}$,…, $x_{q}$, there are only finitely many possibilities for the point $x_{\mathrm{out}}$, depending in a non-trivial way on $x_{1}$,…, $x_{q}$. In particular, for generic $x_{1},\dots,x_{q}$, none of the intersection points of $C$ with $\partial X^{v}$ are intersection of an internal exceptional curve with $\partial X^{v}$, and so $C$ does not contain any internal exceptional curve. If $q=0$ and $x_{\mathrm{out}}$ is an intersection point of an internal exceptional curve $E$ with $\partial X^{v}$, then $\mathcal{O}(C-kE)$ has trivial restriction to $\partial X^{v}$. Since the invariant $N_{\tau}^{\beta}$ is invariant under deformation of $(X^{v},\partial X^{v})$, it follows from [42, Proposition 4.1] that $C=kE$ if $N_{\tau}^{\beta}\neq 0$.
@@
In all cases, we conclude that, if $N_{\tau}^{\beta}\neq 0$, then either $\beta$ is a multiple of an internal exceptional curve, or $\beta$ is represented by a curve in $X^{v}$ containing no internal exceptional curve and no component of $\partial X^{v}$. In the first case, the corresponding walls are given as in Lemma 6.9, and so the finiteness and polynomiality is clear. In the second case, using the notations of Definition 6.8, the class $\beta$ is uniquely determined by $h_{\star}\beta$, and $-F\cdot\beta=-h_{\star}F\cdot h_{\star}\beta$. For a wall non-trivial modulo $J^{k}$, we have $-F\cdot\beta\geq k$, and so $-h_{\star}F\cdot h_{\star}\beta\leq k$. Since $-h_{\star}F$ is $f^{\prime}$-ample by Definition 6.8 of a good divisor, there are finitely many such classes $\beta$. Combined with the induction hypothesis, this implies the finiteness and polynomiality of the walls non-trivial modulo $J^{k}$. ∎
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$, and let $F$ be a good effective divisor on $\mathcal{X}$. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}$, there exists finitely many $\beta\in NE(\mathcal{X}/\overline{\mathcal{X}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$.
@@
First note that, by Lemma 6.7, if $N_{pqr}^{\beta}(\mathcal{X})\neq 0$, then $-F\cdot\beta$ is fixed to be an integer $k$ depending only on $p,q,r$. As reviewed in §5.3, $N_{pqr}^{\beta}(\mathcal{X})$ can be computed in terms of broken lines in $C(P)$, or equivalently in terms of jagged paths in $P$. Moreover, these broken lines have fixed asymptotic directions determined by $p$ and $q$, and fixed final direction determined by $r$, and so the corresponding jagged paths have fixed initial points, given by the radial projections of $p$ and $q$ on $P$, and fixed endpoint, given by the radial projection of $r$ on $P$.
Let $\mathscr{P}^{\prime}$ be the coarsening of $\mathscr{P}$ obtained by removing the edges $e$ such that the corresponding double curve $X^{e}$ of $\mathcal{X}_{0}$ is contained in $G$. Such curves are distinct, and so $\mathscr{P}^{\prime}$ is just obtained by removing diagonals of some quadrilaterals of $\mathscr{P}$. By Equation (5.7), each time a jagged path crosses an edge of $\mathscr{P}^{\prime}$, the corresponding kink is non-trivial modulo $J$, and so the intersection with $-F$ of the curve class attached to the jagged path increases by $1$. Since $-F\cdot\beta=k$ is fixed by $p,q,r$, one deduces that the number of times a jagged path contributing to $N_{pqr}^{\beta}(\mathcal{X})$ crosses an edge of $\mathscr{P}^{\prime}$ is uniformly bounded. Moreover, if $(W,f_{W})$ is a wall generically contained in a 2-dimensional face of $\mathscr{P}^{\prime}$, then $f_{W}=1\mod J$ by Lemma 6.9, and so, by Equation (5.7), the number of times a jagged path crosses such a wall is also uniformly bounded. Hence, we have in fact a uniform bound on the number of times a jagged path can cross a wall. In addition, by Lemma 6.9, there are only finitely many walls non-trivial modulo $J^{k}$, and they all contain only finitely many terms modulo $J^{k}$. Since the jagged paths have fixed asymptotic monomials $z^{p}$ or $z^{q}$, we obtain at every wall-crossing finitely many monomials, and since the number of wall-crossings is uniformly bounded, we conclude that there are finitely many possible jagged paths contributing to structures constants of the form $N_{pqr}^{\beta}(\mathcal{X})$ with fixed $p,q,r$, and so in particular finitely $\beta$ with $N_{pqr}^{\beta}(\mathcal{X})\neq 0$. ∎
@@
    $f:=g\circ h:\mathcal{X}\rightarrow\overline{\mathcal{X}}$ is crepant.
@@
    There exists an effective $f$-nef Cartier divisor $F$ on $\mathcal{X}$, supported on the exceptional locus of $f$, such that $g_{\star}F$ is a Cartier divisor on $\mathcal{Z}$ and $-g_{\star}F$ is $h$-ample.
@@
First recall that by general MMP, [23, Corollary 1.4.3] – see also [59, Corollary 2.6], there exists a crepant $\mathbb{Q}$-factorial terminalization of $\overline{\mathcal{X}}$. Moreover, two such crepant $\mathbb{Q}$-factorial terminalizations are isomorphic in codimension one – see e.g. [61], and so have the same number $e(\overline{\mathcal{X}})$ of exceptional divisors.
We prove the result by induction on $e(\overline{\mathcal{X}})$. If $e(\overline{\mathcal{X}})=0$, then, for any crepant $\mathbb{Q}$-factorial terminalization $f:\mathcal{X}\rightarrow\overline{\mathcal{X}}$, one can trivially take $\mathcal{Z}=\overline{\mathcal{X}}$, $g=f$, $h=\mathrm{Id}$, and $F=0$.
[…]
By the induction hypothesis applied to $\mathcal{Z}^{\prime}$, there exists a crepant $\mathbb{Q}$-factorial terminalization $f^{\prime\prime}=g\circ h^{\prime\prime}:\mathcal{X}\rightarrow\mathcal{Z}^{\prime}$, where $g:\mathcal{X}\rightarrow\mathcal{Z}$ is an isomorphism in codimension one, $h^{\prime\prime}:\mathcal{Z}\rightarrow\mathcal{Z}^{\prime}$, and there exists a $f^{\prime\prime}$-nef Cartier divisor $F^{\prime\prime}$ on $\mathcal{X}$, supported on the exceptional locus of $f^{\prime\prime}$, such that $g_{\star}F^{\prime\prime}$ is a Cartier divisor on $\mathcal{Z}$ and $-g_{\star}F^{\prime\prime}$ is $h^{\prime\prime}$-ample. To prove the result for $\overline{\mathcal{X}}$, we set $f:=f^{\prime\prime}\circ h^{\prime}:\mathcal{X}\rightarrow\overline{\mathcal{X}}$, and $h:=h^{\prime\prime}\circ h^{\prime}:\mathcal{Z}\rightarrow\overline{\mathcal{X}}$. As $-g_{\star}^{\prime}E$ is $h^{\prime}$-ample and $-g_{\star}F^{\prime\prime}$ is $h^{\prime\prime}$-ample, we have that $-g_{\star}F^{\prime\prime}-\lambda g_{\star}^{\prime}E$ is $h$-ample for $\lambda$ large enough, and it is enough to take $F:=(f^{\prime\prime})^{\star}(-g_{\star}F^{\prime\prime}-\lambda g_{\star}^{\prime}E)$ to conclude the proof.

### v3

@@ 6.2. Structure constants under flops
Since any flop can be expressed as a composition of flops of relative Picard rank one, and Theorem 6.7 establishes that flops of relative Picard rank one of a generic quasi-projective open Kulikov degeneration are either multiple M1 or M2 flops, it suffices to prove Theorem 6.14 under the assumption that $b$ is a multiple M1 or M2 flop. For multiple M1 flops, the result follows by combination of Theorem 6.9 and Lemma 6.12. Similarly, for multiple M2 flops, the result follows from Theorem 6.11 in conjunction with Lemma 6.13. ∎
@@ 6.3. The finiteness property
In this section, we prove a finiteness result for the punctured Gromov–Witten invariants $N_{pqr}^{\beta}(\mathcal{X})$, which implies that the mirror algebra of a quasi-projective open Kulikov degeneration $\pi:\mathcal{X}\rightarrow\Delta$ can be defined over $\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$, and not just over the Artinian rings $\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]/I$. We first introduce the notion of a good divisor in Definition 6.16 and prove the finiteness result under the assumption of existence of a good divisor in Theorem 6.19. Then, we show in Lemmas 6.20-6.21 that a good divisor always exist after flops, and we deduce from this the general finiteness result in Theorem 6.24.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. For every divisor $F$ on $\mathcal{X}$ supported on $\mathcal{X}_{0}$, and for every $p,q,r\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$, we have
@@
    The divisor F-F is nef, and for every proper irreducible curve $C\subset\mathcal{X}_{0}$ we have $F\cdot C\geq 1$, except for finitely many disjoint curves $C$ which are either interior exceptional curves, as in Definition 6.1, or (1,1)(-1,-1)-double curves, as in Definition 6.3.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration and $F$ a good effective divisor on $\mathcal{X}$. By Definition 6.16 ii) of a good divisor, the curves $C$ with $F\cdot C=0$ are exactly the curves contracted by a birational morphism, and so they span a face $G$ of the cone $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\otimes\mathbb{R}$. We denote by $J\subset NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ the corresponding monoid ideal consisting of curve classes $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ such that $\beta\notin G$. By Definition 6.16 i) of a good divisor, the curve classes $\beta\in J$ are exactly the curve classes $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ such that $F\cdot\beta\geq 1$.
@@
where $v$ and $v^{\prime}$ are the vertices adjacent to $e$, $(E_{i})_{i\in I_{v,e}}$ (resp. $(E_{j})_{j\in I_{v^{\prime},e}}$) is the set of interior exceptional curves $E_{i}$ in $X^{v}$ (resp . $X^{v^{\prime}}$) intersecting $X^{e}$, and where $m_{v,e}$ (resp.) is the integral primitive direction of $e$ pointing towards $v$ (resp . $v^{\prime}$).
@@
We consider $\mathbb{A}^{1}$-curves $C$ of class $C]\notin J$, that is, $C]\in G$. Since $C$ is connected, and the curves in Definition 6.16 i) spanning $G$ are disjoint and rigid, it follows that $C$ must be either a multiple of an interior exceptional curve or a multiple of a double curve. However, the formal neighborhood of a double curve is toric, whereas the tropicalization of an $\mathbb{A}^{1}$-curve has a unique leg. By the toric balancing condition for punctured curves, this implies that $C$ cannot be a multiple of a double curve. Consequently, $C$ is necessarily a multiple of an interior exceptional curve $E_{i}$, which belongs to an irreducible component $X^{v}$ of $\mathcal{X}_{0}$ and intersects a component $X^{e}\subset X^{v}$ of the double locus of $\mathcal{X}_{0}$. Denote by $f_{W,E_{i}}$ the contribution to $f_{W}$ of the multiple covers of $E_{i}$. The formal neighborhood of $E_{i}$ in the log scheme $\mathcal{X}_{0}^{\dagger}$ is isomorphic to the formal neighborhood of the exceptional curves appearing in the proof of [18, Lemma 4.21]. Therefore, applying [18, Lemma 4.21], we obtain
@@
We prove the result by induction on $k$. For $k=0$, the result follows from Lemma 6.17. For the induction step, let us assume that the result holds for k1k-1, and show that it holds for $k$. For this, first note that every wall $(W,f_{W})$ of $\mathfrak{D}_{\mathcal{X}}$ with $f_{W}\neq 1\mod J^{k}$ and $f_{W}=1\mod J^{k-1}$ has an initial point $x\in P$. By consistency of $\mathfrak{D}_{\mathcal{X}}$, either $x\in P_{\mathbb{Z}}$, or $x$ is an intersection point of walls which are non-trivial modulo $J^{k-1}$. By the induction hypothesis, there are finitely many walls non-trivial modulo $J^{k-1}$, and so finitely many possible points $x$ which are initial points for a wall non-trivial modulo $J^{k}$. Hence, it suffices to prove that for such a point $x$, there are finitely many walls $(W,f_{W})$ with initial point $x$ such that $f_{W}\neq 1\mod J^{k}$, and that for every such a wall, $f_{W}\mod J^{k}$ is a polynomial.
@@
If $x$ is contained in the interior of a 2-dimensional face of $\mathscr{P}$, then $\mathfrak{D}_{\mathcal{X}}$ is given locally around $x$ by a consistent scattering diagram in $\mathbb{R}^{2}$. Moreover, all the walls passing through $x$ are trivial modulo $J$ by Lemma 6.17, and so the walls non-trivial modulo $J^{k}$ are uniquely determined by consistency from the walls non-trivial modulo $J^{k-1}$, and so the finiteness and polynomiality of the walls non-trivial modulo $J^{k}$ follows from the induction hypothesis and [42, Theorem 6.38].
@@
If $x$ is contained in the interior of an edge of $\mathscr{P}$, then $\mathfrak{D}_{\mathcal{X}}$ is still given locally around $x$ by a consistent scattering diagram in $\mathbb{R}^{2}$. The only difference with the previous case is that some walls passing through $x$ might be trivial modulo $J$. However, these walls are of the explicit form given by Lemma 6.17, and so the result still holds by [42, Proposition 6.47].
@@
It remains to consider the case where $x=v$ is a vertex of $\mathscr{P}$, corresponding to an irreducible component $X^{v}$ of $\mathcal{X}_{0}$. It follows from [43, Theorem 8.15] that the punctured Gromov–Witten invariants contributing to the walls $(W,f_{W})$ with initial point $x=v$ and non-trivial modulo $J^{k}$ can be expressed explicitly in terms of punctured Gromov–Witten invariants $N_{\tau}^{\beta}$ of $(X^{v},\partial X_{v})$ and of the walls non-trivial modulo $J^{k-1}$ passing through $v$. More precisely, it follows from [43, Lemma 8.9] that, up to applying corner blow-ups to $(X^{v},\partial X^{v})$, the invariants $N_{\tau}^{\beta}$ are counts of curves $C$ in $X^{v}$ intersecting $\partial X^{v}$ at fixed smooth points $x_{1},\dots,x_{q}$, with multiplicities $m_{1},\dots,m_{q}$, and at a smooth non-fixed point $x_{\mathrm{out}}$ with multiplicity $m_{\mathrm{out}}$. If $q\neq 0$, then, as in the proof of [49, Lemma 4.1], given the points $x_{1}$,…, $x_{q}$, there are only finitely many possibilities for the point $x_{\mathrm{out}}$, depending in a non-trivial way on $x_{1}$,…, $x_{q}$. In particular, for generic $x_{1},\dots,x_{q}$, none of the intersection points of $C$ with $\partial X^{v}$ are intersection of an interior exceptional curve with $\partial X^{v}$, and so $C$ does not contain any interior exceptional curve. If $q=0$ and $x_{\mathrm{out}}$ is an intersection point of an interior exceptional curve $E$ with $\partial X^{v}$, then $\mathcal{O}(C-kE)$ has trivial restriction to $\partial X^{v}$. Since the invariant $N_{\tau}^{\beta}$ is invariant under deformation of $(X^{v},\partial X^{v})$, it follows from [45, Proposition 4.1] that $C=kE$ if $N_{\tau}^{\beta}\neq 0$.
@@
In all cases, we conclude that, if $N_{\tau}^{\beta}\neq 0$, then either $\beta$ is a multiple of an interior exceptional curve, or $\beta$ is represented by a curve in $X^{v}$ containing no interior exceptional curve and no component of $\partial X^{v}$. In the first case, the corresponding walls are given as in Lemma 6.17, and so the finiteness and polynomiality is clear. In the second case, using the notations of Definition 6.16, the class $\beta$ is uniquely determined by $g_{\star}\beta$, and $F\cdot\beta=-g_{\star}F\cdot g_{\star}\beta$. For a wall non-trivial modulo $J^{k}$, we have $F\cdot\beta\leq k$, and so $g_{\star}F\cdot g_{\star}\beta\leq k$. Since $g_{\star}F$ is $h$-ample by Definition 6.16 of a good divisor, there are finitely many such classes $\beta$. Combined with the induction hypothesis, this implies the finiteness and polynomiality of the walls non-trivial modulo $J^{k}$. ∎
We now prove a key finiteness result. Similar results have been previously established in different contexts in [44, Corollary 6.11] and [70, Proposition 15.10].
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$, and let $F$ be a good effective divisor on $\mathcal{X}$. Then, for every $p,q,r\in C(P)_{\mathbb{Z}}$, there exists finitely many curve classes $\beta\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ such that $N_{pqr}^{\beta}(\mathcal{X})\neq 0$.
@@
First note that, by Lemma 6.15, if $N_{pqr}^{\beta}(\mathcal{X})\neq 0$, then $F\cdot\beta$ is fixed to be an integer $k$ depending only on $p,q,r$. As reviewed in §5.3, $N_{pqr}^{\beta}(\mathcal{X})$ can be computed in terms of broken lines in $C(P)$, or equivalently in terms of jagged paths in $P$. Moreover, these broken lines have fixed asymptotic directions determined by $p$ and $q$, and fixed final direction determined by $r$. Therefore, the corresponding jagged paths have fixed initial points, given by the radial projections of $p$ and $q$ on $P$, and fixed endpoint, given by the radial projection of $r$ on $P$.
Let $\mathscr{P}^{\prime}$ be the coarsening of $\mathscr{P}$ obtained by removing the edges $e$ such that the corresponding double curve $X^{e}$ of $\mathcal{X}_{0}$ is contained in $G$. Such curves are distinct, and so $\mathscr{P}^{\prime}$ is just obtained by removing diagonals of some quadrilaterals of $\mathscr{P}$. By Equation (5.7), each time a jagged path crosses an edge of $\mathscr{P}^{\prime}$, the corresponding kink is non-trivial modulo $J$, and so the intersection with F-F of the curve class attached to the jagged path increases by $1$. Since $F\cdot\beta=k$ is fixed by $p,q,r$, one deduces that the number of times a jagged path contributing to $N_{pqr}^{\beta}(\mathcal{X})$ crosses an edge of $\mathscr{P}^{\prime}$ is uniformly bounded. Moreover, if $(W,f_{W})$ is a wall generically contained in a 2-dimensional face of $\mathscr{P}^{\prime}$, then $f_{W}=1\mod J$ by Lemma 6.17, and so, by Equation (5.7), the number of times a jagged path crosses such a wall is also uniformly bounded. Hence, we have in fact a uniform bound on the number of times a jagged path can cross a wall. In addition, by Lemma 6.17, there are only finitely many walls non-trivial modulo $J^{k}$, and they all contain only finitely many terms modulo $J^{k}$. Since the jagged paths have fixed asymptotic monomials $z^{p}$ or $z^{q}$, we obtain at every wall-crossing finitely many monomials. Since the number of wall-crossings is uniformly bounded, we conclude that there exists only finitely many possible jagged paths contributing to structures constants of the form $N_{pqr}^{\beta}(\mathcal{X})$ with fixed $p,q,r$, and so in particular only finitely many curve classes $\beta$ with $N_{pqr}^{\beta}(\mathcal{X})\neq 0$. ∎
@@
    $f:=h\circ g:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is crepant.
@@
    There exists an effective $f$-nef Cartier divisor $F$ on $\mathcal{X}$, supported on the exceptional locus of $f$, such that $g_{\star}F$ is a Cartier divisor on $\mathcal{Z}$ and $g_{\star}F$ is $h$-ample.
@@
First recall that by general MMP, [24, Corollary 1.4.3] – see also [65, Corollary 2.6], there exists a crepant $\mathbb{Q}$-factorial terminalization of $\mathcal{X}^{\mathrm{can}}$. Moreover, any two such crepant $\mathbb{Q}$-factorial terminalizations are isomorphic in codimension one – see e.g. [67], and so have the same number $e(\mathcal{X}^{\mathrm{can}})$ of exceptional divisors.
We prove the result by induction on $e(\mathcal{X}^{\mathrm{can}})$. If $e(\mathcal{X}^{\mathrm{can}})=0$, then, for any crepant $\mathbb{Q}$-factorial terminalization $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$, one can trivially take $\mathcal{Z}=\mathcal{X}^{\mathrm{can}}$, $g=f$, $h=\mathrm{Id}$, and $F=0$. Assume by induction that the result holds for $e(\mathcal{X}^{\mathrm{can}})<n$ with $n\in\mathbb{Z}_{\geq 1}$, and consider $\mathcal{X}^{\mathrm{can}}$ with $e(\mathcal{X}^{\mathrm{can}})=n$. Since $n\geq 1$, by [65, Proposition 2.4] – see also [24, Corollary 1.4.3], there exists a crepant projective birational morphism $f^{\prime}:\mathcal{X}^{\prime}\rightarrow\mathcal{X}^{\mathrm{can}}$, with $\mathcal{X}^{\prime}$ $\mathbb{Q}$-factorial, such that $f^{\prime}$ has exactly one exceptional divisor $E$, and E-E is $f^{\prime}$-nef.
Since $f^{\prime}$ is crepant birational, and E-E is $f^{\prime}$-nef, it follows from the base point free theorem that E-E is $f^{\prime}$-semi-ample. Thus, $f^{\prime}$ factors as $\mathcal{X}^{\prime}\xrightarrow{g^{\prime}}\mathcal{Z}^{\prime}\xrightarrow{h^{\prime}}\mathcal{X}^{\mathrm{can}}$, where $g^{\prime}$ contracts exactly the irreducible curves $C$ with $E\cdot C=0$, and $\mathcal{O}_{\mathcal{X}^{\prime}}(-E)$ is the $g^{\prime}$-pullback of a $h^{\prime}$-ample line bundle on $\mathcal{Z}^{\prime}$. We claim that $g^{\prime}$ is an isomorphism in codimension one. Otherwise, it would contract the unique $f^{\prime}$-exceptional divisor $E$, that is, $g_{\star}^{\prime}(-E)=0$, which is effective. However, since $E$ is $g^{\prime}$-nef and E-E is not effective, this contradicts the negativity lemma [73, Lemma 3.39]. Hence, $g^{\prime}$ is an isomorphism in codimension one, $g_{\star}^{\prime}E$ is Cartier and $h^{\prime}$-ample, and $e(\mathcal{Z}^{\prime})=e(\mathcal{X}^{\mathrm{can}})-1$.
By the induction hypothesis applied to $\mathcal{Z}^{\prime}$, there exists a crepant $\mathbb{Q}$-factorial terminalization $f^{\prime\prime}=h^{\prime\prime}\circ g:\mathcal{X}\rightarrow\mathcal{Z}^{\prime}$, where $g:\mathcal{X}\rightarrow\mathcal{Z}$ is an isomorphism in codimension one, $h^{\prime\prime}:\mathcal{Z}\rightarrow\mathcal{Z}^{\prime}$, and there exists a $f^{\prime\prime}$-nef Cartier divisor $F^{\prime\prime}$ on $\mathcal{X}$, supported on the exceptional locus of $f^{\prime\prime}$, such that $g_{\star}F^{\prime\prime}$ is a Cartier divisor on $\mathcal{Z}$ and $g_{\star}F^{\prime\prime}$ is $h^{\prime\prime}$-ample. To prove the result for $\mathcal{X}^{\mathrm{can}}$, we set $f:=f^{\prime\prime}\circ h^{\prime}:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$, and $h:=h^{\prime}\circ h^{\prime\prime}:\mathcal{Z}\rightarrow\mathcal{X}^{\mathrm{can}}$. As $g_{\star}^{\prime}E$ is $h^{\prime}$-ample and $g_{\star}F^{\prime\prime}$ is $h^{\prime\prime}$-ample, we have that $g_{\star}F^{\prime\prime}-\lambda g_{\star}^{\prime}E$ is $h$-ample for $\lambda$ large enough, and it suffices to take $F:=(f^{\prime\prime})^{\star}(-g_{\star}F^{\prime\prime}-\lambda g_{\star}^{\prime}E)$ to conclude the proof.